# Generating Physically Stable and Buildable Brick Structures from Text

### v2

@@ 5.3 Extensions and Applications
Robotic Assembly of Generated Structures.   We demonstrate automated assembly using a dual-robot-arm system in Figure 8. The robots use the manipulation policy [39] and the asynchronous multi-agent planner [24] to manipulate toy bricks and construct the structure. More details are included in Appendix B. Since the generated structures are physically stable, efficient and automated assembly can be performed.
Manual Assembly.   Our generated brick structures are physically valid and can be assembled by hand. See Appendix C for details.
@@
6 Discussion and Limitations
In this work, we have introduced BrickGPT, an autoregressive model for generating interconnecting toy brick structures from text prompts. Our method learns to predict the next brick sequentially while ensuring physical stability and buildability. We have shown that our method outperforms LLM backbone models and several recent text-to-3D generation methods.
Limitations.   Though our method outperforms existing methods, it still has several limitations. First, due to limited computational resources, we have not explored the largest 3D dataset. As a result, our method is restricted to producing designs within a $20\times 20\times 20$ grid across 21 categories, while recent 3D generation methods can create a wider variety of objects. Future work includes scaling up model training at higher grid resolutions on larger, more diverse datasets, such as Objaverse-XL [7]. Training on large-scale datasets can also improve generalization to out-of-distribution text prompts.
Acknowledgments.   We thank Minchen Li, Ken Goldberg, Nupur Kumari, Ruihan Gao, and Yihao Shi for their discussions and help. We also thank Jiaoyang Li, Philip Huang, and Shobhit Aggarwal for developing the bimanual robotic system. This work is partly supported by the Packard Foundation, Cisco Research Grant, and Amazon Faculty Award. This work is also in part supported by the Manufacturing Futures Institute, Carnegie Mellon University, through a grant from the Richard King Mellon Foundation. KD is supported by the Microsoft Research PhD Fellowship.
@@ 5. Excludes medium-related terms (model, render, design)
StableText2Brick Details.   We generate StableText2Brick starting from the objects in ShapeNetCore [2]. While ShapeNetCore provides both voxel and mesh representations, we find that working directly with mesh data better preserves geometric details. We voxelize the 3D mesh into a $20\times 20\times 20$ grid representation and generate its brick structure using a delete-and-rebuild algorithm. This algorithm is similar to that studied in prior work [46]. However, instead of initializing the structure by filling its voxels with $1\times 1$ unit bricks and randomly merging them into larger bricks, we greedily place bricks to fill the voxels layer-by-layer from bottom to top. We prioritize placing 1) bricks that are only partially supported by bricks on the layer below, 2) bricks that touch multiple bricks on the layer below, 3) large bricks, and 4) bricks of the opposite orientation from bricks on the layer below.
@@
With our delete-and-rebuild mesh-to-brick algorithm, we generate $62{,}000+$ brick structures covering the $21$ categories with $31{,}218$ unique 3D objects from ShapeNetCore. Among the selected 3D objects, $\sim$92(i.e., $28{,}822$) of them have at least one stable brick design, which offer $47{,}000+$ stable layouts. Our StableText2Brick dataset significantly expands upon the previous StableLego dataset [38] in several key aspects: it contains >3×>3$\times$> 3 × more stable unique 3D objects with >5×>5$\times$> 5 × more stable structures compared to $8{,}000+$ in StableLego, spans a diverse set of 21 object categories, and provides detailed geometric descriptions for each shape.
Figure 9: Given a prompt and several example structures in text format, a pre-trained LLaMA model can generate brick designs with some structure.
@@
To force each output brick to be in the format “{h}×{w} ({x},{y},{z})”, we sample only from the set of valid tokens during each step. For example, the first token must be a digit, the second an ‘x’, and so on. This has little to no effect on the output of BrickGPT, but helps force the baseline pre-trained models to output bricks in the specified format.
Only 1.2% of the generated designs exceeded the maximum number of rollbacks and failed to produce a stable final structure.

### v3

@@ 5.3 Extensions and Applications
Robotic Assembly of Generated Structures. We demonstrate automated assembly using a dual-robot-arm system. The robots use the manipulation policy [39] and the asynchronous multi-agent planner [24] to manipulate toy bricks and construct the structure. More details are included in Appendix B. Since the generated structures are physically stable, efficient automated assembly can be performed.
Manual Assembly. Our generated designs are physically valid and can be hand-assembled. See Appendix C.
@@
6 Conclusion
In this work, we have introduced BrickGPT, an autoregressive model for generating interconnecting toy brick structures from text prompts. Our method predicts the next brick sequentially while ensuring physical stability and buildability. We have shown that our method outperforms LLM backbone models and recent text-to-3D generation methods.
Acknowledgments
We thank Minchen Li, Ken Goldberg, Nupur Kumari, Ruihan Gao, and Yihao Shi for their discussions and help. We also thank Jiaoyang Li, Philip Huang, and Shobhit Aggarwal for developing the bimanual robotic system. This work is partly supported by the Packard Foundation, Cisco Research Grant, and Amazon Faculty Award. This work is also in part supported by the Manufacturing Futures Institute, Carnegie Mellon University, through a grant from the Richard King Mellon Foundation. KD is supported by the Microsoft Research PhD Fellowship.
@@ 5. Excludes medium-related terms (model, render, design)
StableText2Brick Details. We generate StableText2Brick starting from the objects in ShapeNetCore [2]. While ShapeNetCore provides both voxel and mesh representations, we find that working directly with mesh data better preserves geometric details. We voxelize the 3D mesh into a $20\times 20\times 20$ grid representation and generate its brick structure using a delete-and-rebuild algorithm. This algorithm is similar to that studied in prior work [46]. However, instead of initializing the structure by filling its voxels with $1\times 1$ unit bricks and randomly merging them into larger bricks, we greedily place bricks to fill the voxels layer-by-layer from bottom to top. We use eight commonly available standard bricks: $1\times 1$, $1\times 2$, $1\times 4$, $1\times 6$, $1\times 8$, $2\times 2$, $2\times 4$, and $2\times 6$. We prioritize placing 1) bricks that are only partially supported by bricks on the layer below, 2) bricks that touch multiple bricks on the layer below, 3) large bricks, and 4) bricks of the opposite orientation from bricks on the layer below.
@@
With our delete-and-rebuild shape-to-brick algorithm, we generate 62,000+62{,}000+ brick structures covering the $21$ categories with 31,21831{,}218 unique 3D objects from ShapeNetCore. Among the selected 3D objects, $\sim$92% (i.e., 28,82228{,}822) of them have at least one stable brick design, which offer 47,000+47{,}000+ stable layouts. Our StableText2Brick dataset significantly expands upon the previous StableLego dataset [38] in several key aspects: it contains $3\times$ more stable unique 3D objects with $5\times$ more stable structures compared to 8,000+8{,}000+ in StableLego, spans a diverse set of 21 object categories, and provides detailed geometric descriptions for each shape.
Figure 8: Few-shot Learning. Given a prompt and several example structures in text format, a pre-trained LLaMA model can generate brick designs with some structure.
@@
To force each output brick to be in the format “{h}×{w} ({x},{y},{z})”, we sample only from the set of valid tokens during each step. For example, the first token must be a digit, the second an ‘×’, and so on. This has little to no effect on the output of BrickGPT, but helps force the baseline pre-trained models to output bricks in the specified format.
In our experiments, only 1.2% of the generated designs exceed the maximum number of rollbacks and fail to produce a stable final structure.
@@ Appendix C Manual Assembly
Appendix D Limitations
Though our method outperforms existing methods, it still has several limitations. First, due to limited computational resources, we have not explored the largest 3D dataset. As a result, our method is restricted to producing designs within a $20\times 20\times 20$ grid across 21 categories, while recent 3D generation methods can create a wider variety of objects. Future work includes scaling up model training at higher grid resolutions on larger, more diverse datasets, such as Objaverse-XL [7]. Training on large-scale datasets can also improve generalization to out-of-distribution text prompts.